# HyperSBINN: A Hypernetwork-Enhanced Systems Biology-Informed Neural Network for Efficient Drug Cardiosafety Assessment

### v1

@@ A.2.1 Late sodium current $I_{\text{Na-L}}$
$I_{\text{Na-L}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{Na-L}}\times g_{\text{Na-L}}(t)\times(V(t)-\mathsf{E}_{\text{Na-L}})\,,$
@@
$g_{\text{Na-L}}(t)$ $=f_{\text{Na-L}}(t)\times h_{\text{Na-L}}(t)$
with these two functions satisfying the following ODEs:
$\frac{{\rm{d}}f_{\text{Na-L}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+30}{5}\right)-f_{\text{Na-L}}(t)}{15}\,,$
$\frac{{\rm{d}}h_{\text{Na-L}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+75.6}{6.3}\right)-h_{\text{Na-L}}(t)}{120+\text{e}^{\frac{V(t)+100}{25}}}\,.$
@@ A.2.2 Sustained transient outward potassium current $I_{\text{TO-S}}$
$I_{\text{TO-S}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{TO-S}}\times g_{\text{TO-S}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,,$
@@
$g_{\text{TO-S}}(t)$ $=f_{\text{TO-S}}(t)\times\left(h_{\text{TO-S}}(t)+\frac{1}{2}\sigma\left(-\frac{V+33.5}{10}\right)\right)$
with these two functions satisfying the following ODEs:
$\frac{{\rm{d}}f_{\text{TO-S}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+3}{15}\right)-f_{\text{TO-S}}(t)}{9\times\sigma\left(-\frac{V(t)+3}{15}\right)+0.5}\,,$
$\frac{{\rm{d}}h_{\text{TO-S}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+33.5}{10}\right)-h_{\text{TO-S}}(t)}{3000\times\sigma\left(-\frac{V(t)+60}{10}\right)+30}\,.$
A.2.3 Slowly activating delayed rectifier potassium Current $I_{\text{K-L}}$
$I_{\text{K-L}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{K-L}}\times g_{\text{K-L}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,,$
@@ A.2.3 Slowly activating delayed rectifier potassium Current $I_{\text{K-L}}$
$g_{\text{K-L}}(t)$ $=f_{\text{K-L}}(t)\times h_{\text{K-L}}(t)$
with these two functions satisfying the following ODEs
$\frac{{\rm{d}}f_{\text{K-L}}}{{\rm{d}}t}(t)$ $=\left(\frac{7.19\cdot 10^{-5}\times(V+30)}{1-\exp(-0.148\cdot(V+30))}+\frac{1.31\cdot 10^{-4}\times(V+30)}{\exp(0.0687\cdot(V+30))-1}\right)$
×$~{}~{}~{}~{}~{}~{}~{}\times\left(\sigma\left(\frac{V(t)-1.5}{16.7}\right)-f_{\text{K-L}}(t)\right)\,,$
$\frac{{\rm{d}}h_{\text{K-L}}}{{\rm{d}}t}(t)$ $=\frac{1}{4}\times\left(\frac{7.19\cdot 10^{-5}\times(V+30)}{1-\exp(-0.148\cdot(V+30))}+\frac{1.31\cdot 10^{-4}\times(V+30)}{\exp(0.0687\cdot(V+30))-1}\right)$
×$~{}~{}~{}~{}~{}~{}~{}\times\left(\sigma\left(\frac{V(t)-1.5}{16.7}\right)-h_{\text{K-L}}(t)\right)\,.$
@@ A.2.4 Inward rectifier current $I_{\text{K-I}}$
$I_{\text{K-I}}(t,\mathsf{X})$ $=\mathsf{G}_{\text{K-I}}\times\left(1+\frac{\mathsf{D}}{\mathsf{IC50}_{\text{K-I}}}\right)^{-1}\ \times g_{\text{K-I}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,,$
where
$g_{\text{K-I}}(t)$ $=\sigma\left(-\frac{V(t)+92}{10}\right)\,.$
@@
$\mathsf{G}_{Kr}$ 0.03 nS
$\mathsf{G}_{Ks}$ 0.1505 nS
$\mathsf{G}_{K1}$ 0.29 nS

### v2

@@ A.2.1 Late sodium current $I_{\text{Na-L}}$
$I_{\text{Na-L}}(t,\mathsf{X})=\mathsf{G}_{\text{Na-L}}\times f_{\text{Na-L}}(t)\times h_{\text{Na-L}}(t)\times(V(t)-\mathsf{E}_{\text{Na-L}})\,$,
@@
$\frac{{\rm{d}}f_{\text{Na-L}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+30}{5}\right)-f_{\text{Na-L}}(t)}{15}\,$, (12)
$\frac{{\rm{d}}h_{\text{Na-L}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+75.6}{6.3}\right)-h_{\text{Na-L}}(t)}{120+\text{e}^{\frac{V(t)+100}{25}}}\,.$ (13)
@@ A.2.2 Sustained transient outward potassium current $I_{\text{To-S}}$
$I_{\text{To-S}}(t,\mathsf{X})=\mathsf{G}_{\text{To-S}}\times f_{\text{To-S}}(t)\times\left(h_{\text{To-S}}(t)+\frac{1}{2}\sigma\left(-\frac{V(t)+33.5}{10}\right)\right)\times(V(t)-\mathsf{E}_{\text{K}})\,$,
@@
$\frac{{\rm{d}}f_{\text{To-S}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(\frac{V(t)+3}{15}\right)-f_{\text{To-S}}(t)}{9\times\sigma\left(-\frac{V(t)+3}{15}\right)+0.5}\,$, (14)
$\frac{{\rm{d}}h_{\text{To-S}}}{{\rm{d}}t}(t)$ $=\frac{\sigma\left(-\frac{V(t)+33.5}{10}\right)-h_{\text{To-S}}(t)}{3000\times\sigma\left(-\frac{V(t)+60}{10}\right)+30}\,.$ (15)
A.2.3 Slowly activating delayed rectifier potassium Current $I_{\text{K-S}}$
$I_{\text{K-S}}(t,\mathsf{X})=\mathsf{G}_{\text{K-S}}\times f_{\text{K-S}}(t)\times h_{\text{K-S}}(t)\times(V(t)-\mathsf{E}_{\text{K}})\,$,
@@ A.2.3 Slowly activating delayed rectifier potassium Current $I_{\text{K-S}}$
$\frac{{\rm{d}}f_{\text{K-S}}}{{\rm{d}}t}(t)$ $=\left(\frac{7.19\cdot 10^{-5}\times(V(t)+30)}{1-\exp(-0.148\cdot(V(t)+30))}+\frac{1.31\cdot 10^{-4}\times(V(t)+30)}{\exp(0.0687\cdot(V(t)+30))-1}\right)$ (16)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \times\left(\sigma\left(\frac{V(t)-1.5}{16.7}\right)-f_{\text{K-S}}(t)\right)\,$,
$\frac{{\rm{d}}h_{\text{K-S}}}{{\rm{d}}t}(t)$ $=\frac{1}{4}\times\left(\frac{7.19\cdot 10^{-5}\times(V(t)+30)}{1-\exp(-0.148\cdot(V(t)+30))}+\frac{1.31\cdot 10^{-4}\times(V(t)+30)}{\exp(0.0687\cdot(V(t)+30))-1}\right)$ (17)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \times\left(\sigma\left(\frac{V(t)-1.5}{16.7}\right)-h_{\text{K-S}}(t)\right)\,.$
@@ A.2.4 Inward rectifier current $I_{\text{K-I}}$
$I_{\text{K-I}}(t,\mathsf{X})=\mathsf{G}_{\text{K-I}}\times\sigma\left(-\frac{V(t)+92}{10}\right)\times(V(t)-\mathsf{E}_{\text{K}})\,.$
@@
$\mathsf{G}_{\text{K-R}}$ 0.03 nS
$\mathsf{G}_{\text{K-S}}$ 0.1505 nS
$\mathsf{G}_{\text{K-I}}$ 0.29 nS
@@ Appendix B ODE Weights $w^{\text{ODE}}$ Initialization Method
Appendix C Additional numerical experiments
C.1 Example of output for two test configurations
Here, we provide comparison between the ODE solver and all proposed alternatives for two drug configurations:
$\mathsf{X}_{1}$ $=\left\{\mathsf{D}=1,\mathsf{IC50}_{\text{K-R}}=12.48,\mathsf{IC50}_{\text{Ca}}=15.7,\mathsf{IC50}_{\text{To-F}}=9.46,\mathsf{IC50}_{\text{Na-F}}=49.4\right\}$ (18)
$\mathsf{X}_{2}$ $=\left\{\mathsf{D}=0.3,\mathsf{IC50}_{\text{K-R}}=0.48,\mathsf{IC50}_{\text{Ca}}=75.7,\mathsf{IC50}_{\text{To-F}}=9.46,\mathsf{IC50}_{\text{Na-F}}=49.4\right\}$ (19)
Figure 5: Comparing the performance for drug configuration $\mathsf{X}_{1}$ (Equation (18) of the proposed model with with the complete loss (HyperSBINN), without the system biology loss (HyperSBINN w/o ODE), without the addition of actual observations (HyperSBINN w/o obs.), or without an hyper network, i.e. with a dense SBINN.
Figure 6: Comparing the performance for drug configuration $\mathsf{X}_{2}$ (Equation (19) of the proposed model with with the complete loss (HyperSBINN), without the system biology loss (HyperSBINN w/o ODE), without the addition of actual observations (HyperSBINN w/o obs.), or without an hyper network, i.e. with a dense SBINN.
Figure 7: Curves of differences between the predicted and true voltage for the hyperSBINN and its version using only observations. One can see that the hyperSBINN predict two curves which performs really bad after $t=400$.